OpenStudy (anonymous):

Any help? Thanks! Cheers! What is the change of entropy for a heat engine using 500.0 J at 20.0°C? 0.04 J/K 0.25 J/K 0.59 J/K 1.71 J/K It takes 3.2 × 103 J of heat to melt an amount of gold (Hf = 6.30 × 104 J/kg). How much gold do you have? 0.051 g 0.51 g 19 g 51 g

OpenStudy (anonymous):

Correct me if I'm wrong but off of the top of my head... (a) \[\Delta S = \frac{ Q }{ T }\] where delta S is the change in entropy, Q is your heat (in Joules) and T is your temperature (in degrees Celsius, add 273 to get to degrees Kelvin)... the calculation should be pretty straightforward, just keep your units in order and you'll get the right answer. (b) E = mC (not Einstein's equation, C will be related to your molar specific heat, or Hf in this question). \[\frac{ 3.2 E103 J }{6.3 E104 J/kg } = answer (\in kilograms)\] as with the first problem keep your units in line and the answer is pretty straightforward.